# Robust Distributed Estimation: Extending Gossip Algorithms to Ranking and Trimmed Means

### v7

@@ 1 Introduction
Distributed learning has gained significant attention in machine learning applications where data is naturally distributed across a network, either due to resource or communication constraints [Chen et al., 2021]. The rapid development of the Internet of Things has further amplified this trend, as the number of connected devices continues to grow, producing large volumes of data at the edge of the network. As a result, edge computing, where computation is performed closer to the data source, has emerged as a viable alternative to conventional cloud computing [Cao et al., 2020, Villar-Rodriguez et al., 2023]. By reducing reliance on centralized servers, edge computing offers several advantages, including lower energy consumption, improved data security, and reduced latency [Cao et al., 2020, Villar-Rodriguez et al., 2023]. In this context, several distributed learning frameworks, including Gossip Learning and Federated Learning, have been developed. Federated Learning depends on a central server for model aggregation [Mammen, 2021], which introduces challenges such as communication bottlenecks and a single point of failure. In contrast, Gossip Learning provides a fully decentralized alternative where nodes communicate only with their nearby neighbors [Hegedűs et al., 2021]. In addition, Gossip learning is particularly well-suited for situations where communication is constrained, such as in peer-to-peer or sensor networks.
@@
One of the central challenges in distributed learning is robustness to corrupted nodes—scenarios where some nodes may contain contaminated data due to hardware faults, or even adversarial attacks [Chen et al., 2017, Yin et al., 2018]. Consider a network of sensors deployed to monitor the temperature in a small region, as described in Boyd et al. [2006]. In idealized scenarios, sensor noise is typically modeled by a zero-mean Gaussian. However, a recent survey (see Ayadi et al. [2017]) mentions that sensor networks are especially prone to outliers due to their reliance on imperfect sensing devices. Indeed, in practice, sensors can malfunction (e.g., become miscalibrated, get stuck at constant values, or report extreme values due to environmental interference). These issues introduce outliers in the dataset which can corrupt the estimated temperature. Specifically, we consider a setting in which a fraction of the nodes hold outlying data [Feng et al., 2014]. In Federated Learning, significant progress has been made in designing robust aggregation methods based on statistics, such as the median and trimmed mean [Blanchard et al., 2017, Baruch et al., 2019, Yin et al., 2018]. These statistics are known to be more robust to outliers than the standard mean [Huber, 1992]. However, ensuring robustness in Gossip Learning remains challenging. Unlike Federated Learning, where a central server can enforce robust aggregation rules, gossip algorithms inherently rely on local communication. To the best of our knowledge, approaches relying on globally estimating robust statistics remain largely unexplored in the gossip literature.
@@
In this paper, we address the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Specifically, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. Ranks play a crucial role in the framework we develop here, as they allow us to identify outliers and compute robust statistics on the one hand Rieder [1982], and because they can be obtained by means of pairwise comparisons on the other. This allows us to exploit the gossip approach proposed in Colin et al. [2015] to compute pairwise averages ($U$-statistics) in a decentralized way. Note that both ranks and trimmed means are inherently global statistics, which makes them more challenging to estimate in a decentralized manner compared to statistics like averages, minima, or maxima that rely on local conservation principles.
@@ 2.1 Problem Formulation and Framework
We consider a decentralized setting where $n\geq 2$ real-valued observations $X_{1},\ldots,X_{n}$ are distributed over a communication network represented by a connected and non-bipartite graph $\mathcal{G}=(V,E)$, see Diestel [2025]: the observation $X_{k}$ is assigned to node $k\in[n]$. For simplicity, we assume no ties: for all $k\neq l$, $X_{k}\neq X_{l}$. Communication between nodes occurs in a stochastic and pairwise manner: at each iteration, an edge of the communication graph $\mathcal{G}$ is chosen uniformly at random, allowing the corresponding neighboring nodes to exchange information. This popular setup is robust to network changes and helps reduce communication overhead and network congestion [Boyd et al., 2006, Duchi et al., 2011, Colin et al., 2015]. We focus on the synchronous gossip setting, where nodes have access to a global clock and synchronize their updates [Boyd et al., 2006, Duchi et al., 2011, Colin et al., 2015]. We assume that a fraction $0<\varepsilon<1/2$ of the data is corrupted and may contain outliers, as stipulated in Huber’s contamination model, refer to [Huber, 1992].
@@
The goal pursued here is to develop a robust gossip algorithm that accurately estimates the mean over the network despite the presence of outliers, specifically the $\alpha$-trimmed mean with $\alpha\in(0,1/2)$, i.e., the average of the middle $(1-2\alpha)$-th fraction of the observations. This statistic, discarding the observations of greater or smaller rank, is a widely used location estimator when data contamination is suspected; see [Van der Vaart, 2000, Serfling, 2009]. Formally, let $X_{n(1)}\leq X_{n(2)}\leq\dots\leq X_{n(n)}$ be the order statistics (i.e., the observations sorted in ascending order), and define $m=\lfloor\alpha n\rfloor$ for $\alpha\in(0,1/2)$. The $\alpha$-trimmed mean is given by:
@@
The framework can be extended to the case of $\ell\geq 2$ tied observations with the mid-rank method: the rank assigned to the $\ell$ tied values is the average of the ranks they would have obtained in absence of ties, that is, the average of $p+1,p+2,\ldots,p+\ell$, which equals $p+(\ell+1)/2$, see Kendall [1945]. To account for the possibility of non-integer rank estimates, one may use the adjusted inclusion interval to define the weights in (2): $I_{n,\alpha}=[u-1/2,v+1/2]$. Note that when ranks are integers, this adjustment has no effect.
@@ 2.2 Related Works – State of the Art
Distributed ranking (or ordering) is considered in Chiuso et al. [2011], where a gossip algorithm for estimating ranks on any communication graph is proposed. The algorithm is proved to converge, in the sense of yielding the correct rank estimate, in finite time with probability one. However, their work does not provide any convergence rate bound, and empirical results suggest that the algorithm is suboptimal in scenarios with long return times, i.e., when the expected time for a random walk on the graph to return to its starting node is long (see the Supplementary Material). Alternatively, here we take advantage of the fact that ranks can be calculated using pairwise comparisons $\mathbb{I}_{\{X_{k}>X_{l}\}}$, so as to build on the GoSta approach in Colin et al. [2015], originally introduced for the distributed estimation of $U$-statistics, with a proved convergence rate bound of order $\mathcal{O}(1/t)$.
@@
To the best of our knowledge, the estimation of $\alpha$-trimmed means has not yet been explored in the gossip literature. Several related works examine the estimation of medians and quantiles in sensor networks [Shrivastava et al., 2004, Greenwald and Khanna, 2004, Kempe et al., 2003]. A key limitation of these works is their reliance on a special node, such as a base station or leader, which initiates queries, broadcasts information, and collects data from other sensors—an assumption that does not apply to our setting [Shrivastava et al., 2004, Greenwald and Khanna, 2004, Kempe et al., 2003]. Another work has proposed an algorithm for estimating quantiles [Haeupler et al., 2018]; however, there seems to be no guarantees of convergence to the true quantiles. Moreover, their algorithm assumes a fully connected communication graph and the ability to sample from four nodes at each step and to sample $K$ random nodes at the end of the protocol, whereas we consider the more challenging setting of arbitrary communication graphs and pairwise communication. Recently, He et al. [2022] proposed a novel local aggregator, ClippedGossip, for Byzantine-robust gossip learning. In our pairwise setup with a fixed clipping radius, their approach—though reasonable for robust optimization—does not work for robust estimation. Specifically, ClippedGossip ultimately converges to the corrupted mean and, therefore, fails to reduce the impact of outliers. A detailed analysis is provided in the Supplementary Material.
@@ 3.1 Algorithm – Convergence Analysis
We introduce GoRank, a gossip algorithm for estimating the ranks of the observations distributed on the network, see Algorithm 1. It builds on GoSta, an algorithm originally designed for estimating pairwise averages ($U$-statistics of degree $2$) proposed and analyzed in Colin et al. [2015]. The GoRank algorithm exploits the fact that ranks can be computed by means of pairwise comparisons:
@@
5:   Update rank estimate: $R_{k}\leftarrow nR^{\prime}_{k}+1$.  // update(k, t)
@@
Let $R_{k}(t)$ and $R^{\prime}_{k}(t)$ denote the local estimates of $r_{k}$ and $r^{\prime}_{k}$ respectively at node $k\in[n]$ and iteration $t\geq 1$. Each node maintains an auxiliary observation, denoted $Y_{k}(t)$, which enables the propagation of observations across the network, despite communication constraints. For each node $k$, the variables are initialized as $Y_{k}(0)=X_{k}$ and $R^{\prime}_{k}(0)=0$. At each iteration $t\geq 1$, node $k$ updates its estimate $R^{\prime}_{k}(t)$ by computing the running average of $R^{\prime}_{k}(t-1)$ and $\mathbb{I}_{\{X_{k}>Y_{k}(t-1)\}}$. The rank estimate is then computed as $R_{k}(t)=nR^{\prime}_{k}(t)+1$. Next, an edge $(i,j)\in E$ is selected uniformly at random, and the corresponding nodes exchange their auxiliary observations: $Y_{i}(t)=Y_{j}(t-1)$ and $Y_{j}(t)=Y_{i}(t-1)$. This random swapping procedure allows each observation to perform a random walk on the graph, which is described by the permutation matrix $\mathbf{W}_{1}(t)=\mathbf{I}_{n}-(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$, which plays a key role in the convergence analysis Colin et al. [2015]. By taking the expectation with respect to the edge sampling process, we obtain $\mathbf{W}_{1}:=\mathbb{E}[\mathbf{W}_{1}(t)]=\mathbf{I}_{n}-({1/|E|})\mathbf{L}$. This stochastic matrix $\mathbf{W}_{1}$ shares similarities with the transition matrix used in gossip averaging Shah et al. [2009], Boyd et al. [2006], Koloskova et al. [2019]: it is symmetric and doubly stochastic. Consequently, $\mathbf{W}_{1}$ has eigenvector $\mathbf{1}_{n}$ with eigenvalue $1$, resulting in a random walk with uniform stationary distribution. In other words, after sufficient iterations (i.e., in a nearly stationary regime), each observation has (approximately) an equal probability of being located at any given node—a property that would naturally hold (exactly) without swapping if the communication graph were complete. In addition, it can be shown that, if the graph is connected and non-bipartite, the spectral gap of $\mathbf{W}_{1}$ satisfies $0<c<1$ and is given by $c=\lambda_{2}/|E|$ where $\lambda_{2}$ is the second smallest eigenvalue (or spectral gap) of the Laplacian Colin et al. [2015]. This spectral gap plays a crucial role in the mixing time of the random walk, reflecting how quickly it (geometrically) converges to its stationary distribution Shah et al. [2009]. In fact, the spectral gap of the Laplacian is also known as the graph’s algebraic connectivity, a larger spectral gap meaning a higher graph connectivity Mohar et al. [1991].
@@
We assume that the network size $n\geq 2$ is known to the nodes. If not, it can be easily estimated by injecting a value +1+1 into the network, where all initial estimates are set to 0, and applying the standard gossip algorithm for averaging Boyd et al. [2006], see the Supplementary Material. The quantities computed for each node will then converge to $1/n$.
@@
We now establish convergence results for GoRank, by adapting the analysis in Colin et al. originally proposed to derive convergence rate bounds for GoSta as follows. For $k\in[n]$, set $\mathbf{h}_{k}=(\mathbb{I}_{\{X_{k}>X_{1}\}},\ldots,\mathbb{I}_{\{X_{k}>X_{n}\}})^{\top}$ and observe that the true rank of observation $X_{k}$ is given by $r_{k}=\mathbf{h}_{k}^{\top}\mathbf{1}_{n}+1$. At iteration $t=1$, the auxiliary observation has not yet been swapped, so the expected estimate is updated as $\mathbb{E}\left[R^{\prime}_{k}(1)\right]=\mathbf{h}_{k}^{\top}\mathbf{e}_{k}$. At the end of the iteration, the auxiliary observation is randomly swapped, yielding the update: $\mathbb{E}[R^{\prime}_{k}(2)]=(1/2)\mathbb{E}[R^{\prime}_{k}(1)]+(1/2)\mathbf{h}_{k}^{\top}\mathbf{W}_{1}\mathbf{e}_{k}$, where $\mathbb{E}[\cdot]$ denotes the expectation taken over the edge sampling process. Using recursion, the evolution of the estimates, for any $t\geq 1$ and $k\in[n]$, is given by
@@
$\mathbb{E}\left[\left|R_{k}(t)-r_{k}\right|\right]\leq\mathcal{O}\left(\frac{1}{\sqrt{ct}}\right)\cdot\sigma_{k}\enspace.$
@@
Building upon the previous analysis, we derive another gossip algorithm called Baseline++, an improved variant of the one proposed by Chiuso et al. for decentralized rank estimation Chiuso et al. [2011]. It is described at length in section A of the Supplementary Material. The algorithm’s steps and propagation closely resemble those of GoRank. However, instead of estimating the means (3) based on pairwise comparisons directly, Baseline++ aims to minimize, for each node $k\in[n]$, a specific ranking loss function, namely the Kendall $\tau$ distance $\phi_{k}(\mathbf{X},\mathbf{R})=\sum_{l=1}^{n}\mathbb{I}_{\left\{\left(X_{k}-X_{l}\right)\cdot\left(R_{k}-R_{l}\right)<0\right\}}$, counting the number of discordant pairs among $((R_{k},X_{k}),\;(R_{l},X_{l}))$ with $l=1,\;\ldots,\;n$. While GoRank provides a quick approximation of the ranks, especially in poorly connected graphs, Baseline++, despite being slower at first, may ultimately achieve a lower overall error as it minimizes a discrete loss function. However, GoRank has a practical advantage: it does not require an initial ranking, which may not always be available in real-world settings. Moreover, it comes with convergence rate guarantees that apply to any graph topology. In contrast, analyzing the convergence of Baseline++ on arbitrary communication graphs is challenging and is left for future work.

### v8

@@ 1 Introduction
Distributed learning has gained significant attention in machine learning applications where data is naturally distributed across a network, either due to resource or communication constraints [chen2021distributed]. The rapid development of the Internet of Things has further amplified this trend, as the number of connected devices continues to grow, producing large volumes of data at the edge of the network. As a result, edge computing, where computation is performed closer to the data source, has emerged as a viable alternative to conventional cloud computing [cao2020overview, villar2023edge]. By reducing reliance on centralized servers, edge computing offers several advantages, including lower energy consumption, improved data security, and reduced latency [cao2020overview, villar2023edge]. In this context, several distributed learning frameworks, including Gossip Learning and Federated Learning, have been developed. Federated Learning depends on a central server for model aggregation [mammen2021federated], which introduces challenges such as communication bottlenecks and a single point of failure. In contrast, Gossip Learning provides a fully decentralized alternative where nodes communicate only with their nearby neighbors [hegedHus2021decentralized]. In addition, Gossip learning is particularly well-suited for situations where communication is constrained, such as in peer-to-peer or sensor networks.
@@
One of the central challenges in distributed learning is robustness to corrupted nodes—scenarios where some nodes may contain contaminated data due to hardware faults, or even adversarial attacks [chen2017distributed, yin2018byzantine]. Consider a network of sensors deployed to monitor the temperature in a small region, as described in [boyd2006randomized]. In idealized scenarios, sensor noise is typically modeled by a zero-mean Gaussian. However, a recent survey (see [ayadi2017outlier]) mentions that sensor networks are especially prone to outliers due to their reliance on imperfect sensing devices. Indeed, in practice, sensors can malfunction (e.g., become miscalibrated, get stuck at constant values, or report extreme values due to environmental interference). These issues introduce outliers in the dataset which can corrupt the estimated temperature. Specifically, we consider a setting in which a fraction of the nodes hold outlying data [feng2014distributed]. In Federated Learning, significant progress has been made in designing robust aggregation methods based on statistics, such as the median and trimmed mean [blanchard2017machine, baruch2019little, yin2018byzantine]. These statistics are known to be more robust to outliers than the standard mean [huber1992robust]. However, ensuring robustness in Gossip Learning remains challenging. Unlike Federated Learning, where a central server can enforce robust aggregation rules, gossip algorithms inherently rely on local communication. To the best of our knowledge, approaches relying on globally estimating robust statistics remain largely unexplored in the gossip literature.
@@
In this paper, we address the problem of robust estimation in gossip algorithms over arbitrary communication graphs. Specifically, we show that an outlier-robust mean can be computed by globally estimating a robust statistic. Ranks play a crucial role in the framework we develop here, as they allow us to identify outliers and compute robust statistics on the one hand [Rieder], and because they can be obtained by means of pairwise comparisons on the other. This allows us to exploit the gossip approach proposed in [colin2015extending] to compute pairwise averages ($U$-statistics) in a decentralized way. Note that both ranks and trimmed means are inherently global statistics, which makes them more challenging to estimate in a decentralized manner compared to statistics like averages, minima, or maxima that rely on local conservation principles.
@@ 2.1 Problem Formulation and Framework
We consider a decentralized setting where $n\geq 2$ real-valued observations $X_{1},\ldots,X_{n}$ are distributed over a communication network represented by a connected and non-bipartite graph $\mathcal{G}=(V,E)$, see [Diestelbook]: the observation $X_{k}$ is assigned to node $k\in[n]$. For simplicity, we assume no ties: for all $k\neq l$, $X_{k}\neq X_{l}$. Communication between nodes occurs in a stochastic and pairwise manner: at each iteration, an edge of the communication graph $\mathcal{G}$ is chosen uniformly at random, allowing the corresponding neighboring nodes to exchange information. This popular setup is robust to network changes and helps reduce communication overhead and network congestion [boyd2006randomized, duchi2011dual, colin2015extending]. We focus on the synchronous gossip setting, where nodes have access to a global clock and synchronize their updates [boyd2006randomized, duchi2011dual, colin2015extending]. We assume that a fraction $0<\varepsilon<1/2$ of the data is corrupted and may contain outliers, as stipulated in Huber’s contamination model, refer to [huber1992robust].
@@
The goal pursued here is to develop a robust gossip algorithm that accurately estimates the mean over the network despite the presence of outliers, specifically the $\alpha$-trimmed mean with $\alpha\in(0,1/2)$, i.e., the average of the middle $(1-2\alpha)$-th fraction of the observations. This statistic, discarding the observations of greater or smaller rank, is a widely used location estimator when data contamination is suspected; see [van2000asymptotic, serfling2009approximation]. Formally, let $X_{n(1)}\leq X_{n(2)}\leq\dots\leq X_{n(n)}$ be the order statistics (i.e., the observations sorted in ascending order), and define $m=\lfloor\alpha n\rfloor$ for $\alpha\in(0,1/2)$. The $\alpha$-trimmed mean is given by:
@@
The framework can be extended to the case of $\ell\geq 2$ tied observations with the mid-rank method: the rank assigned to the $\ell$ tied values is the average of the ranks they would have obtained in absence of ties, that is, the average of $p+1,p+2,\ldots,p+\ell$, which equals $p+(\ell+1)/2$, see [kendall1945treatment]. To account for the possibility of non-integer rank estimates, one may use the adjusted inclusion interval to define the weights in (2): $I_{n,\alpha}=[u-1/2,v+1/2]$. Note that when ranks are integers, this adjustment has no effect.
@@ 2.2 Related Works – State of the Art
Distributed ranking (or ordering) is considered in [chiuso2011gossip], where a gossip algorithm for estimating ranks on any communication graph is proposed. The algorithm is proved to converge, in the sense of yielding the correct rank estimate, in finite time with probability one. However, their work does not provide any convergence rate bound, and empirical results suggest that the algorithm is suboptimal in scenarios with long return times, i.e., when the expected time for a random walk on the graph to return to its starting node is long (see the Supplementary Material). Alternatively, here we take advantage of the fact that ranks can be calculated using pairwise comparisons $\mathbb{I}_{\{X_{k}>X_{l}\}}$, so as to build on the GoSta approach in [colin2015extending], originally introduced for the distributed estimation of $U$-statistics, with a proved convergence rate bound of order $\mathcal{O}(1/t)$.
@@
To the best of our knowledge, the estimation of $\alpha$-trimmed means has not yet been explored in the gossip literature. Several related works examine the estimation of medians and quantiles in sensor networks [shrivastava2004medians, greenwald2004power, kempe2003gossip]. A key limitation of these works is their reliance on a special node, such as a base station or leader, which initiates queries, broadcasts information, and collects data from other sensors—an assumption that does not apply to our setting [shrivastava2004medians, greenwald2004power, kempe2003gossip]. Another work has proposed an algorithm for estimating quantiles [haeupler2018optimal]; however, there seems to be no guarantees of convergence to the true quantiles. Moreover, their algorithm assumes a fully connected communication graph and the ability to sample from four nodes at each step and to sample $K$ random nodes at the end of the protocol, whereas we consider the more challenging setting of arbitrary communication graphs and pairwise communication. Recently, he2022byzantine proposed a novel local aggregator, ClippedGossip, for Byzantine-robust gossip learning. In our pairwise setup with a fixed clipping radius, their approach—though reasonable for robust optimization—does not work for robust estimation. Specifically, ClippedGossip ultimately converges to the corrupted mean and, therefore, fails to reduce the impact of outliers. A detailed analysis is provided in the Supplementary Material.
@@ 3.1 Algorithm – Convergence Analysis
We introduce GoRank, a gossip algorithm for estimating the ranks of the observations distributed on the network, see Algorithm 1. It builds on GoSta, an algorithm originally designed for estimating pairwise averages ($U$-statistics of degree $2$) proposed and analyzed in [colin2015extending]. The GoRank algorithm exploits the fact that ranks can be computed by means of pairwise comparisons:
@@
5:   Update rank estimate: $R_{k}\leftarrow nR^{\prime}_{k}+1$.  // update(k, s)
@@
Let $R_{k}(t)$ and $R^{\prime}_{k}(t)$ denote the local estimates of $r_{k}$ and $r^{\prime}_{k}$ respectively at node $k\in[n]$ and iteration $t\geq 1$. Each node maintains an auxiliary observation, denoted $Y_{k}(t)$, which enables the propagation of observations across the network, despite communication constraints. For each node $k$, the variables are initialized as $Y_{k}(0)=X_{k}$ and $R^{\prime}_{k}(0)=0$. At each iteration $t\geq 1$, node $k$ updates its estimate $R^{\prime}_{k}(t)$ by computing the running average of $R^{\prime}_{k}(t-1)$ and $\mathbb{I}_{\{X_{k}>Y_{k}(t-1)\}}$. The rank estimate is then computed as $R_{k}(t)=nR^{\prime}_{k}(t)+1$. Next, an edge $(i,j)\in E$ is selected uniformly at random, and the corresponding nodes exchange their auxiliary observations: $Y_{i}(t)=Y_{j}(t-1)$ and $Y_{j}(t)=Y_{i}(t-1)$. This random swapping procedure allows each observation to perform a random walk on the graph, which is described by the permutation matrix $\mathbf{W}_{1}(t)=\mathbf{I}_{n}-(\mathbf{e}_{i}-\mathbf{e}_{j})(\mathbf{e}_{i}-\mathbf{e}_{j})^{\top}$, which plays a key role in the convergence analysis [colin2015extending]. By taking the expectation with respect to the edge sampling process, we obtain $\mathbf{W}_{1}:=\mathbb{E}[\mathbf{W}_{1}(t)]=\mathbf{I}_{n}-({1/|E|})\mathbf{L}$. This stochastic matrix $\mathbf{W}_{1}$ shares similarities with the transition matrix used in gossip averaging [shah2009gossip, boyd2006randomized, koloskova2019decentralized]: it is symmetric and doubly stochastic. Consequently, $\mathbf{W}_{1}$ has eigenvector $\mathbf{1}_{n}$ with eigenvalue $1$, resulting in a random walk with uniform stationary distribution. In other words, after sufficient iterations (i.e., in a nearly stationary regime), each observation has (approximately) an equal probability of being located at any given node—a property that would naturally hold (exactly) without swapping if the communication graph were complete. In addition, it can be shown that, if the graph is connected and non-bipartite, the spectral gap of $\mathbf{W}_{1}$ satisfies $0<c<1$ and is given by $c=\lambda_{2}/|E|$ where $\lambda_{2}$ is the second smallest eigenvalue (or spectral gap) of the Laplacian [colin2015extending]. This spectral gap plays a crucial role in the mixing time of the random walk, reflecting how quickly it (geometrically) converges to its stationary distribution [shah2009gossip]. In fact, the spectral gap of the Laplacian is also known as the graph’s algebraic connectivity, a larger spectral gap meaning a higher graph connectivity [mohar1991laplacian].
@@
We assume that the network size $n\geq 2$ is known to the nodes. If not, it can be easily estimated by injecting a value +1+1 into the network, where all initial estimates are set to 0, and applying the standard gossip algorithm for averaging [boyd2006randomized], see the Supplementary Material. The quantities computed for each node will then converge to $1/n$.
@@
We now establish convergence results for GoRank, by adapting the analysis in [colin2015extending], originally proposed to derive convergence rate bounds for GoSta as follows. For $k\in[n]$, set $\mathbf{h}_{k}=(\mathbb{I}_{\{X_{k}>X_{1}\}},\ldots,\mathbb{I}_{\{X_{k}>X_{n}\}})^{\top}$ and observe that the true rank of observation $X_{k}$ is given by $r_{k}=\mathbf{h}_{k}^{\top}\mathbf{1}_{n}+1$. At iteration $t=1$, the auxiliary observation has not yet been swapped, so the expected estimate is updated as $\mathbb{E}\left[R^{\prime}_{k}(1)\right]=\mathbf{h}_{k}^{\top}\mathbf{e}_{k}$. At the end of the iteration, the auxiliary observation is randomly swapped, yielding the update: $\mathbb{E}[R^{\prime}_{k}(2)]=(1/2)\mathbb{E}[R^{\prime}_{k}(1)]+(1/2)\mathbf{h}_{k}^{\top}\mathbf{W}_{1}\mathbf{e}_{k}$, where $\mathbb{E}[\cdot]$ denotes the expectation taken over the edge sampling process. Using recursion, the evolution of the estimates, for any $t\geq 1$ and $k\in[n]$, is given by
@@
$\mathbb{E}\left[\left|R_{k}(t)-r_{k}\right|\right]\leq\mathcal{O}\left(\frac{1}{\sqrt{ct}}\right)\cdot\sigma_{k}\kern 5.0pt.$
@@
Building upon the previous analysis, we derive another gossip algorithm called Baseline++, an improved variant of the one proposed by chiuso2011gossip for decentralized rank estimation [chiuso2011gossip]. It is described at length in section A of the Supplementary Material. The algorithm’s steps and propagation closely resemble those of GoRank. However, instead of estimating the means (3) based on pairwise comparisons directly, Baseline++ aims to minimize, for each node $k\in[n]$, a specific ranking loss function, namely the Kendall $\tau$ distance $\phi_{k}(\mathbf{X},\mathbf{R})=\sum_{l=1}^{n}\mathbb{I}_{\left\{\left(X_{k}-X_{l}\right)\cdot\left(R_{k}-R_{l}\right)<0\right\}}$, counting the number of discordant pairs among $((R_{k},X_{k}),\;(R_{l},X_{l}))$ with $l=1,\;\ldots,\;n$. While GoRank provides a quick approximation of the ranks, especially in poorly connected graphs, Baseline++, despite being slower at first, may ultimately achieve a lower overall error as it minimizes a discrete loss function. However, GoRank has a practical advantage: it does not require an initial ranking, which may not always be available in real-world settings. Moreover, it comes with convergence rate guarantees that apply to any graph topology. In contrast, analyzing the convergence of Baseline++ on arbitrary communication graphs is challenging and is left for future work.